OpenStudy (anonymous):

lim x->infinity cos(pi/x) /-x^-2

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ \cos \frac{ \pi }{ x } }{ x^{-2} }\]

OpenStudy (anonymous):

whoops, -x^-2

OpenStudy (anonymous):